OpenStudy (anonymous):

Can someone please help me? I need to know what the ordered pair is of 2x-6y=36 3x-9y= -27

hero (hero):

y = (36 - 2x)/-6 y = (27+3x)/9 y =y 9(36-2x) = -6(27+3x) 324 - 18x = -162 - 18x no solution

hero (hero):

You don't have any comment on the matter?

OpenStudy (anonymous):

looks correct haha

OpenStudy (anonymous):

Sorry guys I just now saw this post/// Thank you both :D

hero (hero):

It is correct :P